In 2017, Dangerous Dragon, Inc. (a retail clothing company) sold 587,766 units of its product at an average price of $15 per unit. The company reported estimated Returns and allowances in 2017 of 3 percent of gross revenue. Dangerous Dragon actually purchased 550,000 units of its product from its manufacturer in 2017 at an average cost of $9 per unit. Dangerous Dragon began 2017 with 85,000 units of its product in inventory (carried at an average cost of $9 per unit). Operating expenses (excluding depreciation) for Dangerous Dragon, Inc. in 2010 were $ 2,349,110 and depreciation expense was $ 114,372. Dangerous Dragon had $10,000,000 in debt outstanding throughout all of 2010. This debt carried an average interest rate of 6 percent. Finally, Dangerous Dragons tax rate was 40 percent. Dangerous Dragons fiscal year runs from January 1 through December 31. Given this information, construct Dangerous Dragons 2010 multi-step income statement. What did Dangerous Dragon, Inc. record as NET INCOME on its 2010 income statement? Record your answer rounded to the nearest dollar. Do not include a dollar sign or commas in your answer. For example, record $1,234,567.89 as 1234568. For example, record $1,234,567.89 as 1234568 or negative $543,210 as -543210.</font color> Answer: 135,892 x (119,172)Please how to get the answer and explain the steps to get the correct answer listed at the bottom in parenthesis
Answer #1

HI,

For the income Statement we first start with Gross Sales

Gross Sales = number of units sold*per unit price = 587766*15 = $8,816,490

Return = 3% of gross Sales

Hence Net Sales = 8816490*97% = $8,551,995.3

Total Bought units = 550000 at $9

Carry forward inventory= 85000 at $9

but since sold units are 587766 hence we will take cost also fro 587766 units only

So total Cost of good sold = 587766*9 = $5,289,894

So below will be total Income Statement and Net Income for Dangerous Dragon

Income Statement Gross Sales (587766 15) (a) Net Sales d-(a-b) Gross Profit g-d-f 8,816,490.0 Return(3% of gross sales)(b) 264,494.7 8,551,995.3 Cost of Goods Sold(587766* 9)(f) 5,289,894.0 3,262,101.3 Operating Expenses(h) 2,349,110.0 dep. Expenses (i)114,372.0 798,619.3 EBIT(j-g-h-i) Interest Expense(1 mi Profit Before Tax *6%)(k) l-j-k) Tax(40%) (m), 198,619.3 79,447.7 119,171.6 Net Income(I-m)

So Net Income is $119,171.6 = $119,172

Thanks